OpenStudy (anonymous):

find the zeros and the vertex of the equation y=x^2-4x-5

OpenStudy (anonymous):

first coordinate of the vertex is alway \(-\frac{b}{2a}\) which in your case is \(-\frac{-4}{2}=2\) second coordinate of the vertex is what you get for \(y\) when you replace \(x\) by the first coordinate

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

general quadratic looks like \[y=ax^2+bx+c\] you have \[y=x^2-4x-5\]so in your example \[a=1,b=-4,c=-5\]

OpenStudy (anonymous):

the vertex is an ordered pair, looks like \((x, y)\) it is either the highest or lowest point on the parabola. the first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\)

OpenStudy (anonymous):

in your example, \(-\frac{b}{2a}=-\frac{-4}{2\times 1}=2\)

OpenStudy (anonymous):

so x of the vertex is 2?